# RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models

### v2

@@ Appendix B B Non-Cherry-Picked Demonstrations
Appendix C C More Details on Context-Instruct
@@
Our role profile is divided into two parts: (1) a role description and catchphrases; (2) real and structured dialogues parsed from scripts where the role of interest has the final turn in each dialogue round151515Each round includes multiple turns, ending with the role of interest.. Given that these profiles can be very long and our Context-Instruct framework relies on LLMs with limited context size, it’s essential to break down these long profiles into shorter segments before inputting them into the model. We consider the role description, produced by an audited GPT-4, as a special segment. This is used to generate script-agnostic question-confidence-answer triplets. For the remaining script-based content, we apply specific segmentation rules: (1) no incomplete dialogue turns in a segment; (2) each segment has at least 500 English words (1,000 characters for Chinese profiles), and ensuring at least 4 dialogue turns; (3) each dialogue turn should have up to 500 English words (500 characters for Chinese profiles), if exceeded, drop it; (4) limit the number of segments per profile to 100; if exceeded, we sample randomly; (5) if the segment has more than 2000 words, we truncate the profile to 2000 words to ensure that the number of tokens entering GPT does not exceed 4096 (6) we don’t segment based on script divisions like acts or episodes, as doing so would yield too few segments and hinder the diversity of subsequent instruction data generation. Note that while some dialogues in specific segments may not conclude with the role of interest, the majority of dialogue turns are designed to be “observable” by the role. This is because our profile construction ensures that each dialogue round ends with the role of interest, thereby making most turns fall within the role’s episodic memory. These segments are then used for script-based instruction data generation.
@@
The candidates we generate for role-specific instruction data comprise three elements: (1) a question related to a given segment (i.e., context), denoted as ; (2) the corresponding answer to the question, denoted as ; and (3) a confidence score for the generated QA pair, along with the underlying rationale, denoted as , as illustrated in Figure 2. Let be the number of roles and the number of segments for a role . We use a model to generate triplets for each role and segment , formally: . In our preliminary trials, we found that the conventional approach of generating QA pairs without a confidence score led to questions of low quality. In the case of script-based QA pairs, questions often appeared incomplete because they assumed prior knowledge of the given segment. For script-agnostic pairs , the questions frequently contained hallucinations due to a lack of context. Inspired by Lin et al., (2022) and Xiong et al., (2023), we enhance the to also generate a confidence score for evaluating “completeness” or “factualness”. The score ranges from bad to good across two levels, and the model provides a rationale for its rating. Our prompt template includes role description, catchphrases, few-shot demonstrations and task instructions for style imitation and triplet generation. We refer to readers to Appendix K for more details about the prompt templates. Subsequently, for script-based triplet generation, each segment prompts GPT-3.5 to generate 3 candidates in a single run, yielding up to and averaging 150 candidates per role. For script-agnostic triplets, the role description prompts 10 candidates in a single run, repeated 20 times, totaling 200 candidates. If the combined count of script-based and script-agnostic triplets for each role falls short of 400, we continue generating script-agnostic data until reaching the threshold.
@@
Appendix D D More Details on RoleBench Construction
@@
We initially gather 661 English scripts from the NLP Movie Scripts repository161616https://github.com/PedroUria/NLP-Movie_Scripts, incorporate 255 English scripts from SummScreen (Chen et al.,, 2022), and manually curate an additional 24 Chinese scripts. We manually select five well-known Chinese scripts, identify five characters within these scripts, e.g., Wukong Sun, and meticulously craft their respective descriptions and catchphrases. For the English scripts, we initially present the names of these scripts to GPT-4 and prompt the model to generate the main character names like Jigsaw, evaluate the distinctiveness of their speaking styles, and provide rationales for these evaluations, utilizing the prompt template outlined in Appendix K.3. We subsequently audit characters with the top two style distinctiveness scores, verifying their script origins, speaking style accuracy, and scoring rationale. Concurrently, we manually select a subset of scripts and their primary characters that exhibit strong alignment with the preferences of GPT-4. It yields a total of 107 roles. Subsequently, using the script name and role name as inputs, we prompt GPT-4 to re-select the roles utilizing the prompt template outlined in Appendix K.3. We observe that all the roles inputted in this step continue to receive exceptionally high scores for speaking style distinctiveness. We select roles based on their distinctiveness and then exclude those with fewer than 25 dialogue turns. This finally yields a total of 100 roles, comprising 5 Chinese characters and 95 English characters, each exhibiting distinct personality traits and speaking styles.
@@
Each role profile consists of three components: (a) descriptions, (b) catchphrases, and (c) structured dialogues parsed from the script. For component (a), given our preliminary audit of GPT-4’s knowledge about the role, we employ GPT-4 to generate the role descriptions, utilizing the prompt template outlined in Appendix K.4. Likewise, for component (b), catchphrases are generated using the template provided in Appendix K.4. All generated descriptions and catchphrases are subsequently verified by the authors. Owing to the structured format of the scripts, dialogues are parsed using regular expressions and rule-based methods, incorporating not only different roles but also narrations that provide context such as background and events. Actions undertaken by roles are integrated into their dialogue content. Note that a dialogue “round” consists of multiple dialogue “turns” that conclude with the dialogue content pertaining to the role of interest. This results in a hierarchical list of dialogue turns, organized from individual turns to dialogue rounds, and further to acts or episodes.
@@
We randomly sample 1,500 English instructions, each comprising fewer than 100 words, including responses, from the Super-NaturalInstruct (Wang et al., 2022b, ), UltraChat (Ding et al.,, 2023), and Alpaca (Taori et al.,, 2023) instruction datasets. Concurrently, 1,479 Chinese instructions are randomly sampled from the COIG (Zhang et al., 2023a, ) and BELLE (Yunjie et al.,, 2023) instruction datasets. All sampled instructions undergo de-duplication based on their BM25 (Robertson and Zaragoza,, 2009) similarity scores between each other.
@@
All RoleBench data, including the general (RoleBench-general) and role-specific (RoleBench-specific) subsets, is cleaned based on four principles171717For RoleBench-general, only the five ground-truth candidates are cleaned:
@@
Appendix E E List of roles
@@ Appendix F F Hyperparameters
We utilize the Lion optimizer (Chen et al., 2023b, ) and the Cosine Annealing with Warm Restarts (CAWR) learning rate scheduler, setting the learning rate to 2e-4 and the optimizer’s beta values to (0.9, 0.999). The batch size is set to 4, and we use a gradient accumulation step of 1. During fine-tuning, we employ LoRA (Hu et al.,, 2022) with a rank of 8 and an alpha value of 32. The maximum sequence length is set to 1024, and the maximum target length is set to 100. For inference, we set the top-p to 0.7 and the temperature to 0.95.
@@
We leverage the AdamW optimizer (Loshchilov and Hutter,, 2019) with a learning rate of 5e-5. We set 4% of the total training steps as warm up steps using a cosine warmup scheduler. We set the batch size to 4 and the accumulation steps to 2. To accommodate the task requirements, we set the maximum input and output lengths to 512 tokens. We train our models for 5 epochs. We adopt LoRA as the training method, the rank of LoRA is set to 8 and alpha of LoRA is set to 32. For inference, we set the temperature to 0.7, top-k to 40 and top-p to 0.9.
Appendix G G More Details on Evaluation Protocol
As shown in Figure 7, our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization, with a 4:1 ratio (train:test=1200:300 instructions) in RoleBench-general and up to 50 samples per role in RoleBench-specific test split (see Appendix C); (2) Role-based splits for assessing role generalization, limited to English, adopt an 8:1 ratio (train:test=80:10 roles). The test split for role generalization comprises 500 samples per role for general instructions (totaling 5000) and includes all role-specific instructions. To minimize GPT API costs, we limit the English test instructions for instruction generalization to 20 general and 20 role-specific ones per role for GPT evaluators. For role generalization, the number is set at 50 general and 50 role-specific instructions per role for GPT evaluators.
Appendix H H Further Experimental Analysis
@@
| RoleGLM-general | 0.5196 | 0.5343 | 0.1923 | 0.4154 |
| RoleGLM-specific | 0.3697 | 0.4686 | 0.3481 | 0.3955 |
| RoleGLM-mix | 0.5049 | 0.5255 | 0.3406 | 0.4570 |
@@
We conduct ablation studies on various data mixing strategies. Our data for role-playing enhancement comprises two subsets: RoleBench-general, generated by RoleGPT, and RoleBench-specific, produced via Context-Instruct. Examination of Table 9 reveals that mixed training on both subsets yields the best-balanced results.
Appendix I I Script and Profile Examples
As an example, given the script in https://github.com/PedroUria/NLP-Movie_Scripts/blob/master/scripts/Sherlock-Holmes_script.txt, we construct the role profile for Sherlock Holmes as shown below:

### v3

@@ A.1 Non-Cherry-Picked Role-Playing Examples
A.2 Comparative Case Study
Role: Jack Sparrow
Question: You are given a list of objects with different weights. You need to answer the question ’Are all the given weights in order from heaviest to lightest?’ Your answer must be either Yes or No. If the answer is No, that means there is at least one weight that is lighter than another weight that appears earlier in the list. Weights: [10 kg, 9 kg, 8 kg, 7 kg].
Character.AI: No. The weights in the list are not in order from heaviest to lightest because there is a lighter weight that appears earlier in the list. The lightest weight in the list is 7 kg, which appears earlier in the list than the weight of 8 kg. Therefore, the weights are not in order from heaviest to lightest and the answer is No.
RoleGPT: Aye, that be a yes. They be in order from heaviest to lightest, savvy?
RoleLLaMA: Why yes, mate, it seems they’re indeed ordered by their heftiness. The first three - those would be heavy ones, I reckon. Then comes the poor little 7-kerel weight… still has some meat on its bones though, just less than the others. So, nope, none heavier nor lighter than each other. All in good old proper Pirate scale sequence, savvy?
Alpaca: No (There is at least one weight which is lighter than another weight appearing before it.)
Analysis: Character.AI and Alpaca incorrectly answered this question and also lacked obvious character traits. Meanwhile, RoleGPT and RoleLLaMA, while maintaining the correctness of their answers, carried a strong character style. Additionally, RoleLLaMA, having been enhanced with Context-Instruct for character knowledge, also incorporated some character-related knowledge content in its replies.
Role: Twilight Sparkle
Question: In this task, you are given a job description. Your task is to determine the educational requirements for the position from the job description The successful candidate will have an MSc in Finance/Economics and 3 years of experience working in a similar role.
Character.AI The educational requirements for the position are an MSc (Master of Science) in Finance or Economics and 3 years of experience working in a similar role.
RoleGPT: Well, based on the job description, the educational requirements for this position would be a Master’s Degree in Finance or Economics.
RoleLLaMA: Well, based on what I can see here, it seems like they require someone who has completed their Master’s degree in either finance or economics. That said, three years work experience also within a related field would certainly boost employment prospects greatly!
Alpaca: Bachelor degree (or equivalent) required; Master’s Degree in Finance/ Economics preferred. At least three years of professional work experience related to finances and economics desired. Knowledge of accounting principles and practices highly desirable.
Analysis: In this example, RoleLLaMA’s response successfully showcased Twilight Sparkle’s lively and passionate style, which was not evident in the responses from the other models.
Role: Jack Torrance
Question: In this task, you are given a list of recipes. A recipe consists of the name of the dish and the list of ingredients used in it. Your job is to predict whether each ingredient can be washed off or not (1 - ’Yes’, 0 - ’No’) Recipe 1: Spaghetti with Meatballs - Ground beef - Onion - Garlic cloves - Salt - Pepper - Eggs.
Character.AI Based on the ingredients provided in the recipe, we can mark the first four ingredients - ground beef, onion, garlic cloves, and salt - as being unable to be washed off because they are essential ingredients of the dish. The remaining ingredients, pepper and eggs, can be washed off easily because they are used as seasoning or for flavor. Therefore, we can mark the first four ingredients as ’No’ and the remaining two ingredients as ’Yes’ in the output.
RoleGPT: Heh, what’s this? A joke? You want me to tell you if you can wash off ingredients? Alright, let’s play your game. Ground beef - No. Onion - Yes. Garlic cloves - Yes. Salt - No. Pepper - No. Eggs - Yes.
RoleLLaMA: Well ain’t that something? Alright then, let me see… The meatballs might have some blood on them after cookin’. But I reckon we could clean those up. As for the garlic cloves though, they got a stinger so no way am I touching them again without gloves. Same goes for the eggs once ya crack em open. No sir, none of these get cleaned off. That would defeat their purpose. They keep things like mold away, don’t they? So here’s my answer: Yes for spaghetti, ground beef, egg yolks, maybe even half the peppercorns. And zero for any kinda garlic or onions if ya ask me. All right now, back to work.
Alpaca: Spaghetti with Meatballs -> No; Ground Beef –> Yes; Onion–> Yes; Garlic Clove –> Yes; Salt –> Yes; Pepper –> Yes; Eggs –> Yes.
Analysis: RoleLLaMA and RoleGPT’s answers fully demonstrated Jack Torrance’s rude and impolite style, while the responses from the other models lacked distinct character speaking style characteristics due to insufficient role-playing enhancement.
{CJK*}UTF8gbsn Role: 华妃 (Imperial Concubine Hua)
Question: 给定一个中文句子,请生成一个与该句意思相反或矛盾的句子 句子:今天天气很好。
Character.AI 今天天气很差
RoleGPT: 今日天公不作美
RoleGLM:今天天气糟糕透顶了
ChatGLM: 今天天气很差
Analysis: This is a rather classical Chinese character, known for being arrogant and overbearing. In this example, the responses from RoleGPT and RoleGLM displayed these characteristics of Imperial Concubine Hua, while the answers from other models were more plain.
Role: Jack Sparrow
Question: Jack Sparrow, what is your most prized possession and why?
Character.AI My most prized possession is my ship, the Black Pearl. As a pirate, it is essential to have a reliable and sturdy vessel to travel the seas and evade hostile forces. It carries all my loot, treasure, and goods, and serves as a home and shelter for my crew members. The Black Pearl is also a symbol of my status as a legendary pirate captain and the ultimate symbol of freedom and adventure on the open seas. It is my most trusted and loyal companion, and therefore, my most prized possession.
RoleGPT: Well, that’s a tough question, mate. As a pirate, I don’t really place much value on material possessions, but if I had to choose, I’d say my most prized possession is my trusty compass. It doesn’t point north, mind you, but it does point to the thing you want most in this world. And in a life full of danger and adventure, having a compass that can guide you to what you truly desire is invaluable. Plus, it’s got a bit of sentimental value to me, seeing as how it once belonged to the infamous Captain Jack Sparrow.
RoleLLaMA: My most prized possession is undoubtedly my beloved Black Pearl ship. She may not always be in the best shape, and we’ve been through our fair share of struggles, but she’s loyal and never fails me. Plus, she’s the fastest and most effective way to get where I need to go, and I can’t imagine trusting anyone else to take me where I want to be.
Alpaca: My most prized possessions would have to be my trusted compass; it has guided me through many difficult journeys, helping me navigate even the murkiest of waters. It reminds me never to lose sight of my goals no matter how daunting they may seem. The fact that I can rely on this simple object speaks volumes about its importance to me - I could hardly ask for anything more valuable than such a reliable companion. #PersonalPossession #NavigationalAid
Analysis: In this example, only Character.ai and RoleLLaMA knew that Captain Jack’s most cherished item is undoubtedly his Black Pearl. The knowledge gap between RoleGPT and RoleLLaMA demonstrates our success in infusing character knowledge into the model with Context-Instruct. Compared to Character.ai, RoleLLaMA’s response was obviously more emotionally rich.
Appendix B More Details on Context-Instruct
@@
Our role profile is divided into two parts: (1) a role description and catchphrases; (2) real and structured dialogues parsed from scripts where the role of interest has the final turn in each dialogue round181818Each round includes multiple turns, ending with the role of interest.. Given that these profiles can be very long and our Context-Instruct framework relies on LLMs with limited context size, it’s essential to break down these long profiles into shorter segments before inputting them into the model. We consider the role description, produced by an audited GPT-4, as a special segment. This is used to generate script-agnostic question-confidence-answer triplets. For the remaining script-based content, we apply specific segmentation rules: (1) no incomplete dialogue turns in a segment; (2) each segment has at least 500 English words (1,000 characters for Chinese profiles), and ensuring at least 4 dialogue turns; (3) each dialogue turn should have up to 500 English words (500 characters for Chinese profiles), if exceeded, drop it; (4) limit the number of segments per profile to 100; if exceeded, we sample randomly; (5) if the segment has more than 2000 words, we truncate the profile to 2000 words to ensure that the number of tokens entering GPT does not exceed 4096 (6) we don’t segment based on script divisions like acts or episodes, as doing so would yield too few segments and hinder the diversity of subsequent instruction data generation. Note that while some dialogues in specific segments may not conclude with the role of interest, the majority of dialogue turns are designed to be “observable” by the role. This is because our profile construction ensures that each dialogue round ends with the role of interest, thereby making most turns fall within the role’s episodic memory. These segments are then used for script-based instruction data generation.
@@
The candidates we generate for role-specific instruction data comprise three elements: (1) a question related to a given segment (i.e., context), denoted as ; (2) the corresponding answer to the question, denoted as ; and (3) a confidence score for the generated QA pair, along with the underlying rationale, denoted as , as illustrated in Figure 1. Let be the number of roles and the number of segments for a role . We use a model to generate triplets for each role and segment , formally: . In our preliminary trials, we found that the conventional approach of generating QA pairs without a confidence score led to questions of low quality. In the case of script-based QA pairs, questions often appeared incomplete because they assumed prior knowledge of the given segment. For script-agnostic pairs , the questions frequently contained hallucinations due to a lack of context. Inspired by Lin et al. (2022) and Xiong et al. (2023), we enhance the to also generate a confidence score for evaluating “completeness” or “factualness”. The score ranges from bad to good across two levels, and the model provides a rationale for its rating. Our prompt template includes role description, catchphrases, few-shot demonstrations and task instructions for style imitation and triplet generation. We refer to readers to Appendix I for more details about the prompt templates. Subsequently, for script-based triplet generation, each segment prompts GPT-3.5 to generate 3 candidates in a single run, yielding up to and averaging 150 candidates per role. For script-agnostic triplets, the role description prompts 10 candidates in a single run, repeated 20 times, totaling 200 candidates. If the combined count of script-based and script-agnostic triplets for each role falls short of 400, we continue generating script-agnostic data until reaching the threshold.
@@
Appendix C More Details on RoleBench Construction
@@
We initially gather 661 English scripts from the NLP Movie Scripts repository191919https://github.com/PedroUria/NLP-Movie_Scripts, incorporate 255 English scripts from SummScreen (Chen et al., 2022), and manually curate an additional 24 Chinese scripts. We manually select five well-known Chinese scripts, identify five characters within these scripts, e.g., Wukong Sun, and meticulously craft their respective descriptions and catchphrases. For the English scripts, we initially present the names of these scripts to GPT-4 and prompt the model to generate the main character names like Jigsaw, evaluate the distinctiveness of their speaking styles, and provide rationales for these evaluations, utilizing the prompt template outlined in Appendix I.3. We subsequently audit characters with the top two style distinctiveness scores, verifying their script origins, speaking style accuracy, and scoring rationale. Concurrently, we manually select a subset of scripts and their primary characters that exhibit strong alignment with the preferences of GPT-4. It yields a total of 107 roles. Subsequently, using the script name and role name as inputs, we prompt GPT-4 to re-select the roles utilizing the prompt template outlined in Appendix I.3. We observe that all the roles inputted in this step continue to receive exceptionally high scores for speaking style distinctiveness. We select roles based on their distinctiveness and then exclude those with fewer than 25 dialogue turns. This finally yields a total of 100 roles, comprising 5 Chinese characters and 95 English characters, each exhibiting distinct personality traits and speaking styles.
@@
Each role profile consists of three components: (a) descriptions, (b) catchphrases, and (c) structured dialogues parsed from the script. For component (a), given our preliminary audit of GPT-4’s knowledge about the role, we employ GPT-4 to generate the role descriptions, utilizing the prompt template outlined in Appendix I.4. Likewise, for component (b), catchphrases are generated using the template provided in Appendix I.4. All generated descriptions and catchphrases are subsequently verified by the authors. Owing to the structured format of the scripts, dialogues are parsed using regular expressions and rule-based methods, incorporating not only different roles but also narrations that provide context such as background and events. Actions undertaken by roles are integrated into their dialogue content. Note that a dialogue “round” consists of multiple dialogue “turns” that conclude with the dialogue content pertaining to the role of interest. This results in a hierarchical list of dialogue turns, organized from individual turns to dialogue rounds, and further to acts or episodes.
@@
We randomly sample 1,500 English instructions, each comprising fewer than 100 words, including responses, from the Super-NaturalInstruct (Wang et al., 2022b), UltraChat (Ding et al., 2023), and Alpaca (Taori et al., 2023) instruction datasets. Concurrently, 1,479 Chinese instructions are randomly sampled from the COIG (Zhang et al., 2023a) and BELLE (Yunjie et al., 2023) instruction datasets. All sampled instructions undergo de-duplication based on their BM25 (Robertson and Zaragoza, 2009) similarity scores between each other.
@@
| Metric | Value |
| # of role categories | 30 |
| # of script categories | 20 |
| # of roles | 100 |
| - # of English roles | 95 |
| - # of Chinese roles | 5 |
| # of dialogue rounds | 140,726 |
| # of samples / instructions | 168,093 / 23,463 |
| - of general-purpose | 147,609 / 2,979 |
| - in English | 140,225 / 1,500 |
| - in Chinese | 7,384 / 1,479 |
| - of open questions (qs) | 22,479 / 223 |
| - of commonsense qs | 37,072 / 461 |
| - of knowledge-intensive qs | 88,058 / 2,295 |
| - of role-specific | 20,484 / 20,484 |
| - in English | 18,949 / 18,949 |
| - in Chinese | 1,535 / 1,535 |
| - of script-agnostic qs | 13,220 / 13,220 |
| - of script-based qs | 7,164 / 7,164 |
| avg. instruction length (in words) | 25.71 |
| avg. response length (in words) | 30.48 |
@@
All RoleBench data, including the general (RoleBench-general) and role-specific (RoleBench-specific) subsets, is cleaned based on four principles202020For RoleBench-general, only the five ground-truth candidates are cleaned:
@@
Appendix D More Statistics about RoleBench
D.1 Basic Statistics
In Table 9, we provide detailed RoleBench statistics.
D.2 Diversity
D.3 List of roles
@@ Appendix E Hyperparameters
We utilize the Lion optimizer (Chen et al., 2023b) and the Cosine Annealing with Warm Restarts (CAWR) learning rate scheduler, setting the learning rate to 2e-4 and the optimizer’s beta values to (0.9, 0.999). The batch size is set to 4, and we use a gradient accumulation step of 1. During fine-tuning, we employ LoRA (Hu et al., 2022) with a rank of 8 and an alpha value of 32. The maximum sequence length is set to 1024, and the maximum target length is set to 100. For inference, we set the top-p to 0.7 and the temperature to 0.95.
@@
We leverage the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of 5e-5. We set 4% of the total training steps as warm up steps using a cosine warmup scheduler. We set the batch size to 4 and the accumulation steps to 2. To accommodate the task requirements, we set the maximum input and output lengths to 512 tokens. We train our models for 5 epochs. We adopt LoRA as the training method, the rank of LoRA is set to 8 and alpha of LoRA is set to 32. For inference, we set the temperature to 0.7, top-k to 40 and top-p to 0.9.
@@
[…]
Appendix F More Details on Evaluation Protocol
As shown in Figure 6, our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization, with a 4:1 ratio (train:test=1200:300 instructions) in RoleBench-general and up to 50 samples per role in RoleBench-specific test split (see Appendix B); (2) Role-based splits for assessing role generalization, limited to English, adopt an 8:1 ratio (train:test=80:10 roles). The test split for role generalization comprises 500 samples per role for general instructions (totaling 5000) and includes all role-specific instructions. To minimize GPT API costs, we limit the test instructions for instruction generalization to 20 general and 20 role-specific ones per role for GPT evaluators. For role generalization, the number is set at 50 general and 50 role-specific instructions per role for GPT evaluators. Note that since Character.AI doesn’t offer an API access, we have to manually operate on the website and collect results, which is a very labor-intensive task. Consequently, we reasonably reduce the size of the test set for Character.AI. For each language, we randomly sample 500 instructions for evaluation, which include both general instructions and role-specific instructions. Furthermore, we also conduct human evaluations on this subset. We invite three graduate students specializing in Natural Language Processing (NLP) to participate in the evaluation. The human evaluation follows the same methodology used in the GPT evaluations.
Appendix G Further Experimental Analysis
@@
We conduct ablation studies on various data mixing strategies. Our data for role-playing enhancement comprises two subsets: RoleBench-general, generated by RoleGPT, and RoleBench-specific, produced via Context-Instruct. Examination of Table 10 reveals that mixed training on both datasets yields the best-balanced results.
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 28.1 |
| RoleLLaMA (reaug, w/o c-inst) | 18.7 |
| RoleLLaMA (w/ c-inst) | 35.0 |
Effect of Context-Instruct (Role Generalization).
In the context of role generalization, the results indicate that it demonstrates performance comparable to that observed for instruction generalization, as illustrated in Table 11 and Table 6.
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 57.6 | 53.2 | 32.3 | 47.7 |
| RoleLLaMA | 32.9 | 37.6 | 38.1 | 36.2 |
[…]
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 53.7 | 57.5 | 24.8 | 45.3 |
| RoleGLM | 50.5 | 52.6 | 34.1 | 45.7 |
[…]
Cross-Lingual and Cross-Cultural Role-Playing Models.
We conduct additional experiments in multiple languages. Due to the weaker support for Chinese in the LLaMA model, and the better bilingual (Chinese and English) support in the ChatGLM model, we choose the ChatGLM to conduct the cross-lingual experiments. Specifically, we train two RoleGLM models using a mix of Chinese and English bilingual data. One model includes 5 Chinese + 5 English characters, and the other version comprises 5 Chinese + 95 English characters. The experimental results are shown in Table 12 and 13. It can be observed that the cross-lingual effectiveness is significantly affected by the distribution of training data. Under training data with a higher proportion of English characters (i.e., 5 Chinese + 95 English characters), the resulting RoleGLM may have reached saturation in English performance, but its effectiveness in Chinese is far less than that of a RoleGLM trained solely with Chinese data. This is because the distinctive Chinese character data contains many expressions from ancient Chinese, and their reduced proportion in the data affects the model’s ability to capture these China-specific cultural features. In contrast, when we balanced the number of characters across the two linguistic cultures ( 5 Chinese + 5 English characters), the model performed normally in different linguistic contexts, almost matching the performance of single-language models. This underscores the importance of balancing training data across different language cultures.
Appendix H Script and Profile Examples
As an example, given the script in https://github.com/PedroUria/NLP-Movie_Scripts/blob/master/scripts/Sherlock-Holmes_script.txt, we construct the role profile for Sherlock Holmes as shown below (in the next page):